OpenStudy (anonymous):

Please help!!! Formula is w1 • d1= w2 • d2. Solve for d

OpenStudy (anonymous):

can you write the formula again please its kind of confusing

OpenStudy (anonymous):

Is it look clear now?

OpenStudy (anonymous):

w^1 * d^2 = w^2 *d^2

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@inkyvoyd

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

I think d = 0

OpenStudy (anonymous):

I want to how do I solve for d.

OpenStudy (campbell_st):

well if you divide both sides of the equation by w^2 you get 1/w * d^1 = d^2 now divide both sides by d 1/w = d hope it helps and makes sense.

OpenStudy (anonymous):

but we have two w which w^1 and w^2

OpenStudy (anonymous):

how do I eliminate that first?

OpenStudy (anonymous):

ok sorry, thanks alot @campbell_st ! :) I saw how you did.

OpenStudy (anonymous):

Solve for d:d^2 w = d^2 w^2 Subtract d^2 w^2 from both sides d^2 (w-w^2) = 0 Divide both sides by w-w^2 d^2 = 0 Take the square root of both sides Answer: d = 0

OpenStudy (anonymous):

the variable to solve for @Loser66

OpenStudy (anonymous):

so you are telling me to divide with all d with d?

OpenStudy (anonymous):

distance

OpenStudy (anonymous):

weights and distances

OpenStudy (anonymous):

Is that what you are asking?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Actually I am the one who is stupids. :( it isnt you. I am really struggling to understand it. I am not getting the process

OpenStudy (anonymous):

9th

OpenStudy (anonymous):

literal equation

OpenStudy (anonymous):

isnt 9th grade albegra 2?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

I think I understand bit better than earlier way you explained @severus4ever and @campbell_st. Thank you guys helping me so much!!!!!!!!!!! You guys are truly really great!!!!!